Determine whether or not the triangles are similar 30 50° Similar Not similar
Answer #1

A E 30 56 С B D an ABC LA+CBtc(=180 90+so+L(= 180 Lo= 180- 140 Lo= 40 A DEF an ABC ABC and (hoven) LA= LD-90° LE LE Both tria